OpenStudy (anonymous):

Using a generating function, how many ways can you make change for a dollar using pennies, nickels, dimes, quarters, and half-dollars? so far i got (1+x+x^2+x^3+...+x^100) <- pennies (1+x^5+x^10+...+x^100) <- nickels (1+x^10+x^20+...+x^100) <- dimes (1+x^25+x^50+x^75+x^100) <- quarters (1+x^50+x^100) <- half dollars and the generating function is: 1/(1-x)(1-x^5)(1-x^10)(1-x^25)(1-x^50) what I need help is what do I do from here to get the answer I need? Please explain each step you did

OpenStudy (tkhunny):

Hmmm... I like your original FINITE polynomials, but I'm not quite sure how you managed that Generating Function. 1/(1-x) does not result in a finite series.

OpenStudy (anonymous):

the pennies can be written as 1/(x-1)

OpenStudy (tkhunny):

Rethink. \[1 + x + x^2 + ... + x^{100} = \frac{1-x^{101}}{1-x}\]

OpenStudy (anonymous):

http://aleph.math.louisville.edu/teaching/2009FA-681/notes-090917.pdf question 3 and answer 3

OpenStudy (tkhunny):

Please read more carefully. It is the case that \[1 + x + x^{2} + ... + x^{100} = \frac{1 - x^{101}}{1-x}\] as I have stated it. It is also the case that \[1 + x + x^{2} + ... = \frac{1}{1-x}\] It is also the case that in some cases it is simpler to work with the infinite form, rather than the more cumbersome finite form.

OpenStudy (anonymous):

well if i had it right earlier then can you tell me how to do the next step then?

OpenStudy (tkhunny):

If it were me, I'd use that result you posted in the link and divide out the dollar factor. I am not aware of any easy way to pick out one coefficient amongst the hundreds presented in this complicated expression.

OpenStudy (anonymous):

thats the problem i am having what i really wanted to know is how to multiply this out to find the coefficient of x^100 (1+x+x^2+x^3+...+x^100)(1+x^5+x^10+...+x^100)(1+x^10+x^20+...+x^100)(1+x^25+x^50+x^75+x^100)(1+x^50+x^100) multiplying this out would give me a 500th degree polynomial and the answer would be the coefficient of x^100 which i dont know how to get

OpenStudy (tkhunny):

Yeah, you don't want to do that. Try this: http://www.cut-the-knot.org/ctk/GeneratingFunctions.shtml It is not a simple task at all.

OpenStudy (tkhunny):

Of course, using my previuos idea of dividing from the 293 answer given in your original link, it should be immedaitely obvious that using dollars adds only one way to get $1.00. Thus, 292.

OpenStudy (anonymous):

yea that is the answer, but i wanted to show my work instead of just putting the answer down and thanks for the link i take a look at it